Assume that the one-year interest rate in Singapore is 2 per cent. The one-year Australian interest rate is 4 per cent. The spot rate of the Singapore dollar (S$) is A$0.9776. The forward rate of the Singapore dollar is A$1.0465.

Quoted Price

a. Is covered interest arbitrage feasible for Australian investors? Show the results if an Australian company engages in covered interest arbitrage to support your answer.
b. Assume that the spot rate and interest rates remain unchanged as coverage interest arbitrage is attempted by Australian investors. Do you think the forward rate of the Singapore dollar will be affected? If so, state whether it will increase or decrease, and explain why.
Answer #1
a]
1] Forward premium on the S$ = 1.04564/0.9776-1 = 6.96%
2] Difference in interest rates = 4%-2% = 2.00%
3] As the forward premium is not equal to the
difference in interest rates, CIA is possible.
Since the interest rate is less than the forward
premium, it would be profitable to borrow in the
currency having higher interest rate and then
investing in the currency having lower interest
rate.
So borrowing should be in A$ and investing in S$.
Assuming that borrowings are made for A$1000.
The maturity value of the borrowing would be
1000*1.04 = A$1040.
The A$1000 would be converted to S$ at spot to
get 1000/0.9776 = S$1022.91
This would be invested at 2% to have a maturity
value of 1022.91*1.02 = S$1043.37
Simultaneously a forward contract to sell
S$1043.37 at 1.0465 would be entered into.
After 1 year the investment would be realized
and converted to A$ to get 1043.37*1.0465 =
S$1091.89
With these A$ the borrowings having a maturity
value of A$1040 would be repaid and the profit
would be 1091.89-1040 = A$51.89.
b] Yes the forward rate of S$ would be affected as
more and more arbitrageurs would engage in
similar profiting.
The result would be a greater supply of S$ in the
forward market resulting in the decrease of the
forward rate for the Singapore dollar.